OpenStudy (kamille):

Hello, I know it is easy, but can someone explain me how to solve for x? (It is a system of equation, I dont know how to make bracket sign here) \[\frac{ xy }{ 2 }=24\] \[(\frac{ x }{ 2 })^{2}=16\]

OpenStudy (anonymous):

in the first one, you would multiply by two so you would have xy=48, then divide by x and get x=48/y

OpenStudy (anonymous):

the second one you would square root first to get x/2=4 then you would multiply by 2 and get x=8 i hope these help!

OpenStudy (kamille):

oh, that really stupid haha :D i have thought i need to insert x somewhere and get y ! Thanks a lot for your help!

OpenStudy (anonymous):

careful here

OpenStudy (anonymous):

\[(\frac{x}{2})^2=16\] \[\frac{x}{2}=\pm4\]

OpenStudy (anonymous):

oh that is true @satellite73 ! sorry, i forgot that if you introduce the square root it is both positive and negative, good catch!

OpenStudy (kamille):

well, it can't be -4, because it is problem of geometry (i didnt write a problem, because I knew how to make a system of eq), but thanks @satellite73 too!